What is 15% of 97?
What percent of 160 is 70?
200% of what is 90?
Find 93% of 157.
Show all work for each problem.

Answer by TimothyLamb(4379)

0.15*97 = 14.55
---
70/160*100 = 43.75%
---
2.00x = 90
x = 90/2
x = 45
---
0.93*157 = 146.01
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php